using the unit circle o solve for the following must show full work
cot -41pi
----
6

Answer by stanbon(75887)

cot [-41pi/6]
-----
(-41/6) = -6 5/6
------
Comment:: -6 pi is just 3 times around the circle.
----
The reference angle is 1/6 pi
-5/6 pi is the same as +7/6 pi
-----
cot(-41pi/6) = cot(7pi/6) = cot(pi/6) = sqrt(3)
--------------
Cheers,
Stan H.
================